A metallic rod is being heated in an oven where the temperature, T, varies with time, t,
as follows: T = t/5 + 100. The length of the rod, L, varies with temperature, and therefore with time according to the formula: L = 100 + t 10000. Find L as a function of T.
1
Expert's answer
2020-06-29T14:04:57-0400
"T=\frac{t}{5}+100\to t=5T-500"

"L = 100 +10000t=100 +10000(5T-500)\\L=100 +50000(T-100)"